Consider the decisions of Virgin Galactic as it opens an exclusive line between Earth and the Moon. If Virgin Galactic is the sole airline operating this line, it acts as a monopolist in this market. The demand for flights to the Moon is segmented into two groups. Business travelers, who often have to buy last minute tickets, have a fully inelastic demand of 30 for any price between zero and 120 and zero demand for any price above 120. Families going on vacation, who can plan ahead, have a demand given by P = 120 − Q. Although Richard Branson (the founder of Virgin Galactic) had to pay billions of pounds to build the Virgin Galactic ship, by using renewable energy he cut his marginal cost to zero.


a) Assume first that Virgin Galactic is allowed to charge different prices to these different groups (for example, it can charge a different price if you purchase the ticket a few months in advance than if you purchase the ticket last minute). Calculate the monopoly prices charged for both business travelers __________ and families________ , and the total consumer surplus in the whole market____________.


The government is concerned about the monopoly power of Virgin Galactic. To measure the "social cost" of the Virgin Galactic pricing strategy you calculated in (a), the government compares the total surplus you found in (a) to the total surplus that would arise if Virgin Galactic would charge prices as if it was in a competitive equilibrium.


b) As the first step to help the government's calculation, compute the total demand for flights in the market_____________ .


c) With the total demand you have calculated above, calculate the price__________ , the producer surplus__________ , and the consumer surplus__________ that would arise if this was a perfectly competitive market.


d) To get the government measure of the "social cost" of Virgin Galactic pricing strategy, calculate the difference between the total surplus in the perfectly competitive market, and the total surplus you have computed in (a).________


1
Expert's answer
2021-01-25T13:50:11-0500

a) Let's find the monopoly prices charged.

The price for business travelers is P = 120.

The price for families is:

MR = MC,

MR = TR'(Q) = 120 - 2Q,

120 - 2Q = 0,

Q = 60 units,

P = 120 - 60 = 60.

The total consumer surplus in the whole market is "CS = 0 + 0.5\times60\times60 = 1800."

b) The total demand for flights in the market is 120 - Q = 0, Q = 120 units.

c) With the total demand calculated above, the price is P = 120 - 120 = 0, the producer surplus is 0, and the consumer surplus is "CS = 0.5\times120\times120 = 7200", that would arise if this was a perfectly competitive market.

d) The difference between the total surplus in the perfectly competitive market, and the total surplus computed in (a) is 7200 - 1800 = 5400.